Help me please, and show work thanks.
Find the next three terms in the geometric sequence.
1. 3,15,75,375...
thanks show work if some.

Answer by richard1234(7193)

Common ratio is 15/3 = 5
375*5 = 1875
1875*5 = 9375
9375*5 = 46875